OpenStudy (anonymous):

prove the identity cos4θ = 1- 8 sin^2θ cos^2θ

OpenStudy (***[isuru]***):

first use \[\cos 4\theta = 1 - 2 \sin^{2}2\theta \]

OpenStudy (***[isuru]***):

then use \[\sin 2 \theta = 2\sin \theta \cos \theta\] to expand \[\sin^{2 }2\theta\]

OpenStudy (***[isuru]***):

then u will get ur answer :) ...can u do it ?